# Generative Bayesian optimization: generative models as acquisition functions

### v1

@@ 6 Conclusion
This work introduces Generative Bayesian Optimization (GenBO), a unifying framework that turns any generative model into a sampler whose density tracks Bayesian-optimization acquisition functions or preference losses. By eliminating intermediate regression or classification surrogates, GenBO reduces approximation error, simplifies the pipeline to learning just a single generative model, and scales naturally to large batches and high-dimensional or combinatorial design spaces. Theoretical results show convergence to the global optimum, and experiments on sequence and protein design tasks demonstrate competitive performance with more complex surrogate-guided baselines. A few challenges remain. For some variants, GenBO requires choosing and fixing the prior before optimization, and its performance depends on sensible settings of utility and temperature parameters. Despite these caveats, GenBO’s minimal moving parts and principled acquisition-driven training mark a simpler and more scalable alternative to multi-stage guided generation methods.
Appendix
Appendix A Learning parametric models with RKHS convex losses
@@
The assumption above allows us to consider functions $g_{*}$ which cannot be perfectly approximated by the model, though which yet live in the same underlying Hilbert space ${\mathcal{{H}}}_{k}$. The reproducing kernel assumption is also mild, as it simply means that function evaluations are continuous (i.e., well behaved), which can usually not be guaranteed in other types of Hilbert spaces, such as, e.g., ${\mathcal{{L}}}_{2}$-spaces. In fact, every Hilbert space of functions where evaluation functionals are continuous is an RKHS by definition (Steinwart & Christmann, 2008, Def. 4.18). Lastly, we note that we can always find a RKHS that contains the models, such as the minimal construction below.
@@
This is a direct application of classic RKHS results (e.g., Steinwart & Christmann, 2008, Thm. 4.21) where we are treating $\phi:{\boldsymbol{\mathbf{x}}}\mapsto g({\boldsymbol{\mathbf{x}}},\cdot)$ as a feature map mapping into an existing Hilbert space ${\mathcal{{H}}}_{\Theta}$ and taking advantage of its structure to define a new one. ∎
@@
If the RKHS in 1 is insufficiently small to contain the function $g_{*}$ of interest, we can always combine two RKHS to produce a third one containing all elements of the two. Namely, if $g_{*}\in{\mathcal{{H}}}_{*}\neq{\mathcal{{H}}}_{g}$ with kernel $k_{*}:{\mathcal{{X}}}\times{\mathcal{{X}}}\to\mathbb{{R}}$, we can define $k:=k_{*}+k_{g}$, so that ${\mathcal{{H}}}_{k}:={\mathcal{{H}}}_{*}\oplus{\mathcal{{H}}}_{g}$ is also a RKHS (Steinwart & Christmann, 2008; Saitoh & Sawano, 2016).
@@
$\forall s\in\mathbb{{R}},\quad\mathbb{E}[\exp(s\epsilon_{t})\mid{\mathfrak{{F}}}_{t-1}]\leq\exp\left(\frac{s^{2}\sigma_{\epsilon}^{2}}{2}\right).$ (33)
@@
Note that most loss functions in the deep learning literature satisfy the assumptions above, including the squared error and the cross entropy loss. The original Bradley-Terry model in the DPO paper (Rafailov et al., 2023) is not strongly convex, whereas its robust version (Chowdhury et al., 2024), which accounts for preference noise, can be shown to satisfy strong convexity and smoothness.
@@
In this case, a quadratic regularization term $\lVert{\theta}\rVert_{2}^{2}$ corresponds to the choice of a linear kernel, i.e., $k_{\Theta}({\theta},{\theta}^{\prime})={\theta}\cdot{\theta}^{\prime}$, which might appear quite restrictive, as it assumes that our models are linear functions of the parameters. However, note that, for overparameterized neural networks, at the infinite-width limit the model is actually linear in the parameters (Jacot et al., 2018). If we want to be more parsimonious, alternatively, we can choose $k_{\Theta}$ as a universal kernel, such as the squared-exponential, yet preferably not translation invariant, so that $k_{\Theta}({\theta},{\theta})$ is not a constant. One kernel satisfying such assumption would be the exponential dot-product kernel $k_{\Theta}({\theta},{\theta}^{\prime}):=\exp({\theta}\cdot{\theta}^{\prime})$, which is universal for continuous functions over compact subsets of $\Theta$. Nevertheless, we do not impose restrictions on the form of the regularization term $\bar{R}_{n}$, except for the one below, which is followed by our assumptions on the loss.
@@
Lemma 2 (Abbasi-Yadkori, 2012, Cor. 3.6).
@@
Substituting the upper bounds in equations 54 and 56 into Equation 53 and applying the concentration inequality in Equation 48 yields:
@@
which also holds uniformly over all $n\in\mathbb{{N}}$ with probability at least $1-\delta$. Combining Equation 58 and 50 leads to the final result, which concludes the proof. ∎
Despite the model being potentially non-linear and the loss not being required to be least-squares, Theorem 1 shows that we recover the same kind of RKHS-based error bound found in the kernelized bandits literature (Chowdhury & Gopalan, 2017; Durand et al., 2018; Oliveira et al., 2021), up to an approximation gap $\xi_{k}$ w.r.t. the true function $g_{*}$. If identifiability holds, we have $\xi_{k}=0$, and we recover the usual bounds (Durand et al., 2018). [^2] If we further assume that the model can represent any $g\in{\mathcal{{H}}}_{k}$, the factor of 2 multiplying $\beta_{n}$ would also disappear, as the extra $\beta_{n}$ arises from a bound over $\lVert g_{{\theta}_{n}}-g_{n}\rVert$, which would vanish. Alternatively, in the case of neural networks, we can increase the width of the network over time (making sure the model scales up with the data is not uncommon in deep learning approaches), which would then lead to the model covering a whole RKHS, determined by the NTK (Jacot et al., 2018). In general, for a rich enough model class, one may expect $\xi_{k}$ to be small.
@@
Lemma 3 (GP variance upper bound (Steinberg et al., 2025, Lem. E.5)).
@@
Appendix B Additional Experimental Detail
B.1 Text optimization
We use the same annealing threshold scheme for setting $\tau_{t}$ as Steinberg et al. (2025, Eqn. 20), where we set $\eta$ such that when begin at $p_{0}=0.5$ we end at $p_{T}=0.99$. For the proposal distribution, we found these short sequences we best generated by the simple mean-field categorical model,
@@ B.1 Text optimization
where $x_{m}\in{\mathcal{{V}}}$ and $\phi_{m}\in\mathbb{{R}}^{|{\mathcal{{V}}}|}$, and we directly optimise $\phi$. VSD and CbAS use the simple MLP classifier guide in Figure 3.
@@
B.2 Protein design
We use the same threshold function and setting for all of the protein design experiments as in subsection B.1. However, these tasks require a more sophisticated generative model that can capture local and global relationships that relate to protein’s 3D structure. For this we use the auto-regressive (causal) transformer architecture also used in Steinberg et al. (2025),
@@ B.2 Protein design
See for the latter see Phuong & Hutter (2022, Algorithm 10 & Algorithm 14) for maximum likelihood training and sampling implementation details respectively. We give the architectural configuration for the transformers in each task in Table 1, and the classifier CNN used by VSD and CbAS is in Figure 3.
@@ B.3 GenBO settings
sEI Soft Expected Improvement
@@
MF Mean-field proposal model
@@
frequent-reg More frequent regularization (change in $\lambda_{n}$ schedule rate)
reg0p1000 Base regularization factor set to $\lambda_{0}:=0.1$
@@
noprior No informative prior, i.e., $p_{0}({\boldsymbol{\mathbf{x}}})\propto 1$
prior Pre-trained prior, learned from initial (randomly initialized) data ${\mathcal{{D}}}_{0}$
logits Importance weights
@@
Table 3 presents our settings for the different GenBO variants across experiments. The settings for our proposal models followed VSD’s configurations. Our regularization scheme penalized the Euclidean distance between the model’s parameters and their random initialization (He et al., 2015) with $R_{n}({\theta}):=\lambda_{n}\lVert{\theta}-{\theta}_{0}\rVert_{2}^{2}$, using an annealed regularization factor $\lambda_{n}:=\frac{1}{n}\lambda_{0}\log^{2}n$, similar to Dai et al. (2022), which ensures enough exploration, while still $\lambda_{n}\to 0$ as $n\to\infty$, allowing for convergence to the optimal $\theta$$_{*}$. For threshold-based utilities, we set $\tau$ to follow an annealing schedule ranging from the 50% (i.e., the median) to the 99% percentile of the observations marginal distribution for both GenBO and VSD.

### v2

@@ 7 Conclusion
This work introduces Generative Bayesian Optimization (GenBO), a unifying framework that turns any generative model into a sampler whose density tracks BO acquisition functions. We have shown that loss functions over generative models, such as DPO and KL divergences, can be applied to directly learn samplers for batch BO. By eliminating intermediate regression or classification surrogates, GenBO reduces approximation error, simplifies the pipeline to learning just a single generative model, and scales naturally to large batches and high-dimensional or combinatorial design spaces. Theoretical results show convergence to the target distribution, and experiments on text optimization and protein design tasks demonstrate competitive performance with more complex surrogate-guided baselines. A few challenges remain. For some variants, GenBO requires choosing and fixing the prior before optimization, and its performance depends on sensible settings of utility and temperature parameters, whose theory could be further explored. Another avenue is the adaptation to acquisition strategies not expressible as expected utilities, such as Thompson sampling and upper confidence bound. Despite these caveats, GenBO’s minimal moving parts and principled acquisition-driven training mark a simpler and more scalable alternative to multi-stage guided generation methods.
Appendix A Gaussian processes for BO
Assume a Gaussian process prior over $f$, e.g., $f\sim\mathcal{GP}(0,k)$, where $k:{\mathcal{{X}}}\times{\mathcal{{X}}}\to\mathbb{{R}}$ is a positive-definite kernel (Rasmussen2006). Then, given a set of observations ${\mathcal{{D}}}_{n}:=\{{\boldsymbol{\mathbf{x}}}_{i},y_{i}\}_{i=1}^{n}$, corrupted by Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma_{\epsilon}^{2})$, the posterior $f|{\mathcal{{D}}}_{n}\sim\mathcal{GP}(\mu_{n},k_{n})$ is available in closed form with mean and covariance function given by:
$\mu_{n}({\boldsymbol{\mathbf{x}}})$ $:={\boldsymbol{\mathbf{k}}}_{n}({\boldsymbol{\mathbf{x}}})^{\mathsf{T}}({\boldsymbol{\mathbf{{K}}}}_{n}+\sigma_{\epsilon}^{2}{\boldsymbol{\mathbf{{I}}}})^{-1}{\boldsymbol{\mathbf{y}}}_{n}$ (26)
$k_{n}({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}^{\prime})$ $:=k({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}^{\prime})-{\boldsymbol{\mathbf{k}}}_{n}({\boldsymbol{\mathbf{x}}})^{\mathsf{T}}({\boldsymbol{\mathbf{{K}}}}_{n}+\sigma_{\epsilon}^{2}{\boldsymbol{\mathbf{{I}}}})^{-1}{\boldsymbol{\mathbf{k}}}_{n}({\boldsymbol{\mathbf{x}}}^{\prime})$ (27)
$\sigma_{n}^{2}({\boldsymbol{\mathbf{x}}})$ $:=k_{n}({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}),$ (28)
where ${\boldsymbol{\mathbf{k}}}_{n}({\boldsymbol{\mathbf{x}}}):=[k({\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}_{i})]_{i=1}^{n}\in\mathbb{{R}}^{n}$, ${\boldsymbol{\mathbf{{K}}}}_{n}:=[k({\boldsymbol{\mathbf{x}}}_{i},{\boldsymbol{\mathbf{x}}}_{j})]_{i,j=1}^{n}\in\mathbb{{R}}^{n\times n}$, ${\boldsymbol{\mathbf{y}}}_{n}:=[y_{i}]_{i=1}^{n}\in\mathbb{{R}}^{n}$, for ${\boldsymbol{\mathbf{x}}},{\boldsymbol{\mathbf{x}}}^{\prime}\in{\mathcal{{X}}}$. With these closed-form expressions, GP models allow BO algorithms to quantify uncertainty and assess expected utilities of their decisions. However, note that, due to matrix inversions, exact GP inference incurs a computational cost of ${\mathcal{{O}}}(n^{3})$. Hence, one often has to resort to low-rank approximations to make GP predictions tractable in cases involving large amounts of data, such as batch evaluations with large batch size. Alternatively, one may completely discard the GP models and use other surrogates, such as neural networks, and there has been an increasing literature on how to reliably quantify uncertainty for BO when using these models (Li2024).
Appendix B The GenBO algorithm
Input: Domain $\mathcal{{X}}$}, initial data ${\mathcal{{D}}}_{0}$
for $t\in\{1,\dots,T\}$ do
 $q_{t}\in\operatornamewithlimits{argmin}_{q\in{\mathcal{{Q}}}}L_{t-1}(q)$
 // Fit proposal distribution
 ${\mathcal{{B}}}_{t}\overset{i.i.d.}{\sim}q_{t}$
 // Sample batch
 $y_{t,i}\leftarrow f({\boldsymbol{\mathbf{x}}}_{t,i})+\epsilon_{t,i}$, for $i\in\{1,\dots,B\}$
 // Collect observations
 ${\mathcal{{D}}}_{t}={\mathcal{{D}}}_{t-1}\cup\{{\boldsymbol{\mathbf{x}}}_{t,i},y_{t,i}\}_{i=1}^{B}$
 // Update data
Algorithm 1 GenBO
Appendix C Learning parametric models with RKHS convex losses
@@
The assumption above allows us to consider functions $g_{*}$ which cannot be perfectly approximated by the model, though which yet live in the same underlying Hilbert space ${\mathcal{{H}}}_{k}$. The reproducing kernel assumption is also mild, as it simply means that function evaluations are continuous (i.e., well behaved), which can usually not be guaranteed in other types of Hilbert spaces, such as, e.g., ${\mathcal{{L}}}_{2}$-spaces. In fact, every Hilbert space of functions where evaluation functionals are continuous is an RKHS by definition (Steinwart2008, Def. 4.18). Lastly, we note that we can always find a RKHS that contains the models, such as the minimal construction below.
@@
This is a direct application of classic RKHS results (e.g., Steinwart2008, Thm. 4.21) where we are treating $\phi:{\boldsymbol{\mathbf{x}}}\mapsto g({\boldsymbol{\mathbf{x}}},\cdot)$ as a feature map mapping into an existing Hilbert space ${\mathcal{{H}}}_{\Theta}$ and taking advantage of its structure to define a new one. ∎
@@
If the RKHS in 1 is insufficiently small to contain the function $g_{*}$ of interest, we can always combine two RKHS to produce a third one containing all elements of the two. Namely, if $g_{*}\in{\mathcal{{H}}}_{*}\neq{\mathcal{{H}}}_{g}$ with kernel $k_{*}:{\mathcal{{X}}}\times{\mathcal{{X}}}\to\mathbb{{R}}$, we can define $k:=k_{*}+k_{g}$, so that ${\mathcal{{H}}}_{k}:={\mathcal{{H}}}_{*}\oplus{\mathcal{{H}}}_{g}$ is also a RKHS (Steinwart2008; Saitoh2016).
@@
$\forall s\in\mathbb{{R}},\quad\mathbb{E}[\exp(s\epsilon_{t})\mid{\mathfrak{{F}}}_{t-1}]\leq\exp\left(\frac{s^{2}\sigma_{\epsilon}^{2}}{2}\right),\quad\forall t\in\mathbb{{N}}\,.$ (34)
@@
Note that most loss functions in the deep learning literature satisfy the assumptions above, including the squared error and the cross entropy loss. The original Bradley-Terry model in the DPO paper (Rafailov2023dpo) is not strongly convex, whereas its robust version (Chowdhury2024rdpo), which accounts for preference noise, can be shown to satisfy strong convexity and smoothness.
@@
In this case, a quadratic regularization term $\lVert{\theta}\rVert_{2}^{2}$ corresponds to the choice of a linear kernel, i.e., $k_{\Theta}({\theta},{\theta}^{\prime})={\theta}\cdot{\theta}^{\prime}$, which might appear quite restrictive, as it assumes that our models are linear functions of the parameters. However, note that, for overparameterized neural networks, at the infinite-width limit the model is actually linear in the parameters (Jacot2018). If we want to be more parsimonious, alternatively, we can choose $k_{\Theta}$ as a universal kernel, such as the squared-exponential, yet preferably not translation invariant, so that $k_{\Theta}({\theta},{\theta})$ is not a constant. One kernel satisfying such assumption would be the exponential dot-product kernel $k_{\Theta}({\theta},{\theta}^{\prime}):=\exp({\theta}\cdot{\theta}^{\prime})$, which is universal for continuous functions over compact subsets of $\Theta$. Nevertheless, we do not impose restrictions on the form of the regularization term $\bar{R}_{n}$, except for the one below, which is followed by our assumptions on the loss.
@@
Lemma 2 (Abbasi-Yadkori2012, Cor. 3.6).
@@
Substituting the upper bounds in equations 55 and 57 into Equation 54 and applying the concentration inequality in Equation 49 yields:
@@
which also holds uniformly over all $n\in\mathbb{{N}}$ with probability at least $1-\delta$. Combining Equation 59 and 51 leads to the final result, which concludes the proof. ∎
Despite the model being potentially non-linear and the loss not being required to be least-squares, Theorem 1 shows that we recover the same kind of RKHS-based error bound found in the kernelized bandits literature (Chowdhury2017; Durand2018; Oliveira2021), up to an approximation gap $\xi_{k}$ w.r.t. the true function $g_{*}$. If identifiability holds, we have $\xi_{k}=0$, and we recover the usual bounds (Durand2018). [^2] If we further assume that the model can represent any $g\in{\mathcal{{H}}}_{k}$, the factor of 2 multiplying $\beta_{n}$ would also disappear, as the extra $\beta_{n}$ arises from a bound over $\lVert g_{{\theta}_{n}}-g_{n}\rVert$, which would vanish. Alternatively, in the case of neural networks, we can increase the width of the network over time (making sure the model scales up with the data is not uncommon in deep learning approaches), which would then lead to the model covering a whole RKHS, determined by the NTK (Jacot2018). In general, for a rich enough model class, one may expect $\xi_{k}$ to be small.
@@
Lemma 3 (GP variance upper bound (Steinberg2025variational, Lem. E.5)).
@@
Appendix D Additional Experimental Detail
D.1 Text optimization
We use the same annealing threshold scheme for setting $\tau_{t}$ as Steinberg2025variational, where we set $\eta$ such that when begin at $p_{0}=0.5$ we end at $p_{T}=0.99$. For the proposal distribution, we found these short sequences we best generated by the simple mean-field categorical model,
@@ D.1 Text optimization
where $x_{m}\in{\mathcal{{V}}}$ and $\phi_{m}\in\mathbb{{R}}^{|{\mathcal{{V}}}|}$, and we directly optimize $\phi$. VSD and CbAS use the simple MLP classifier guide in Figure 4.
@@
D.2 Protein design
We use the same threshold function and setting for all of the protein design experiments as in subsection D.1. However, these tasks require a more sophisticated generative model that can capture local and global relationships that relate to protein’s 3D structure. For this we use the auto-regressive (causal) transformer architecture also used in Steinberg2025variational,
@@ D.2 Protein design
See for the latter see phuong2022formal for maximum likelihood training and sampling implementation details respectively. We give the architectural configuration for the transformers in each task in Table 1, and the classifier CNN used by VSD and CbAS is in Figure 4.
@@ D.3 GenBO settings
sEI Soft Expected Improvement, i.e., $\mathrm{softplus}(y-\tau)$
@@
MF Mean-field categorical proposal model
@@
fr More frequent regularization (change in $\lambda_{n}$ schedule rate)
r0p10 Base regularization factor set to $\lambda_{0}:=0.1$
@@
np No (informative) prior, i.e., $p_{0}({\boldsymbol{\mathbf{x}}})\propto 1$
p Pre-trained prior, learned from initial (randomly initialized) data ${\mathcal{{D}}}_{0}$
lg Importance weights
lr0p10 Learning rate setting for training the generative model (e.g., 0.1 in this case)
pcmin0p50 Minimum percentile for threshold $\tau_{t}$ annealing schedule (e.g., 50% in this case)
pcmax0p90 Maximum percentile for threshold $\tau_{t}$ annealing schedule (e.g., 90% in this case)
@@
Table 3 presents our settings for the different GenBO variants across experiments. The settings for our proposal models followed VSD’s configurations. Our regularization scheme penalized the Euclidean distance between the model’s parameters and their random initialization (He2015nninit) with $R_{n}({\theta}):=\lambda_{n}\lVert{\theta}-{\theta}_{0}\rVert_{2}^{2}$, using an annealed regularization factor $\lambda_{n}:=\frac{1}{n}\lambda_{0}\log^{2}n$, similar to Dai2022, which ensures enough exploration, while still $\lambda_{n}\to 0$ as $n\to\infty$, allowing for convergence to the optimal $\theta$$_{*}$. For threshold-based utilities, we mainly set the quantile threshold $\tau_{t}$ to follow an annealing schedule ranging from the 50% (i.e., the median) to the 99% percentile of the observations marginal distribution for both GenBO and VSD, where the percentile $\gamma_{t}$ corresponding the quantile is updated as $\gamma_{t}:=\gamma_{t-1}^{\eta}$, where $\eta:=\left(\frac{\log\gamma_{T}}{\log\gamma_{0}}\right)^{\frac{1}{T-1}}\in(0,1)$.
D.4 Results summary
Besides the plots in section 6, we summarize the final results in Table 4 and 5.
ALOHA Ehrlich-15 Ehrlich-32 Ehrlich-64
Random mut. 3.80 $\pm$ 0.40
LaMBO-2 0.19 $\pm$ 0.17 0.36 $\pm$ 0.15 0.95 $\pm$ 0.02
CbAS 2.20 $\pm$ 0.40 0.57 $\pm$ 0.12 0.61 $\pm$ 0.10 0.98 $\pm$ 0.01
GA 0.45 $\pm$ 0.12 0.61 $\pm$ 0.10 0.98 $\pm$ 0.01
VSD 0.00 $\pm$ 0.00 0.19 $\pm$ 0.17 0.32 $\pm$ 0.09 0.97 $\pm$ 0.00
GenBO 0.20 $\pm$ 0.40 0.00 $\pm$ 0.00 0.28 $\pm$ 0.16 0.94 $\pm$ 0.02
Table 4: Final average regret (lower is better) for the best-performing variant of each method across the ALOHA (text optimization) and Ehrlich benchmarks
FoldX (Stability) FoldX (SASA)
Random mut. 2.79 $\pm$ 0.22 12550.26 $\pm$ 56.34
LaMBO-2 3.19 $\pm$ 0.58 12456.10 $\pm$ 126.64
CbAS 3.65 $\pm$ 0.23 12376.65 $\pm$ 298.30
VSD 4.20 $\pm$ 0.42 12537.97 $\pm$ 186.35
GenBO 3.28 $\pm$ 0.35 13285.42 $\pm$ 221.60
Table 5: FoldX average maximum outcome for the best-performing variant of each method
Method Average Runtime
CbAS 53.38 s $\pm$ 2.05 s
VSD 42.89 s $\pm$ 2.56 s
GenBO 14.88 s $\pm$ 0.26 s
Table 6: Summary of average run times with standard deviations on the ALOHA text optimization problem.
D.5 Ablations
This section presents ablation studies. We vary the minimum and maximum percentile for the threshold annealing settings of both GenBO (with PI utility) and VSD on the text optimization problem in Figure 5. In Figure 6, we vary the batch size $B$ for GenBO on the Ehrlich benchmark problem of sequence length 32.
(a) GenBO
(b) VSD
Figure 5: Final average simple regret for GenBO and VSD as a function of the minimum and maximum percentile in the annealing schedule.
Figure 6: Batch evaluation size $B$ ablation on Ehrlich benchmark of length 32. The plot presents the final average simple regret for each $B$ setting, with error bars corresponding to $\pm 1$ standard deviation. All variants were run for the same number of BO rounds as in the original experiment.